If one oscillation has 4.4 times the energy of a second one of equal frequency and mass, what is the ratio of their amplitudes?

Express your answer using two significant figures.

Ahighenergy/Alowenergy
Answer #1

E = kA^2/2
Let k = 2nd spring constant.
(k)A1^2/2 = 4.4(kA2^2/2) ==> A1/A2 = sqrt(9/1)=3:1